OpenStudy (anonymous):

What is the ratio of the circumferences for two circles with areas 6π m2 and 150π m2? 1:10 1:25 1:5 1:50

OpenStudy (anonymous):

area =pi*r^2=pi*6 given Area =pi*R^2=pi*150 given circum.=2*pi*r=2*pi*sq.rt6 circum.=2*pi*R^2=2*pi*sq.rt150 ratio =1/5